OpenStudy (anonymous):

Jaren is the manager of a department store. The morning of a big holiday sale, when Jaren opened the door, 12 customers were waiting outside to come into the store. Each hour the store is open, another 13 customers enter the store, and 8 customers leave the store. Use this information to complete the table below. Hours Open Number of Customers In the Store 0 12 1 17 2 22 3 4 The x-values represent the number of hours the store is open. The y-values represent the number of customers in the store. The equation which represents the number of customers in the store is

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Each hour the store is open, another 13 customers enter the store, and 8 customers leave the store.

OpenStudy (igreen):

What's 13 - 8?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, so 5 people enter the store every hour.

OpenStudy (igreen):

So add 5 to 22.

OpenStudy (igreen):

To find out how many people are in the store on the 3rd hour.

OpenStudy (anonymous):

27

OpenStudy (igreen):

Yes, now add 5 to that again to find out how many people were in the store on the 4th hour.

OpenStudy (anonymous):

@iGreen The first second and third are already done :/

OpenStudy (anonymous):

And 32 BTW

OpenStudy (igreen):

Oh..so where are you?

OpenStudy (anonymous):

On the 3 and 4 hours open

OpenStudy (anonymous):

Hello?@iGreen

OpenStudy (igreen):

We just found those out.. 3rd hour = 27 4th hour = 32

OpenStudy (igreen):

Is that it?

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

But i have more

OpenStudy (anonymous):

The equation which represents the number of customers in the store is y _= x +_ .

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Is it 'y = x + _'

OpenStudy (igreen):

What we were adding to each of those numbers back there?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

The x-values represent the number of hours the store is open. The y-values represent the number of customers in the store.

OpenStudy (igreen):

Yes, so that's what goes in the blank. y = x + 5

OpenStudy (anonymous):

Heres something to help us..

OpenStudy (igreen):

That answers all your questions.

OpenStudy (anonymous):

So what we put in the chart are the answers?

OpenStudy (anonymous):

@iGreen .Or do we add all of the people then thats y and we add the hours to get x?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@phi

OpenStudy (igreen):

???

OpenStudy (anonymous):

@iGreen Finnally

OpenStudy (igreen):

Yeah..

OpenStudy (igreen):

27 and 32

OpenStudy (igreen):

Hours Open Number of Customers In the Store 0 12 1 17 2 22 3 27 4 32

OpenStudy (anonymous):

Well look at my last question

OpenStudy (anonymous):

Do i add the hours for X?And do i add the people for y?

OpenStudy (igreen):

I don't understand what you're trying to say.. As 'x' increases by 1, 'y' increases by 5.

OpenStudy (anonymous):

The x-values represent the number of hours the store is open. The y-values represent the number of customers in the store. The equation which represents the number of customers in the store is y = _x + _.

OpenStudy (phi):

I think they want the equation of the line that goes through your "points" pick any two entries, say (0,12) and (1,17) and find the equation

OpenStudy (igreen):

Oh

OpenStudy (igreen):

Find the slope and plug it into point-slope form.

OpenStudy (igreen):

Along with one of the points.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

\(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{17-12}{1-0}\)

OpenStudy (igreen):

Simplify

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

17 - 12 = ? 1 - 0 = ? Divide after

OpenStudy (anonymous):

5 divided by zero?

OpenStudy (igreen):

1 - 0 isn't 0

OpenStudy (anonymous):

Ohhhh yeah 1 xD sorry

OpenStudy (igreen):

Yep, so 5 divided by 1.

OpenStudy (anonymous):

5

OpenStudy (igreen):

Yes, that is the slope, so this is what we have so far: y = 5x + _

OpenStudy (igreen):

The next one is the y-intercept, which is what 'y' is when x = 0..we can find this on our table..

OpenStudy (anonymous):

Then what would y be?

OpenStudy (anonymous):

Oh so y=0 because x isnt?

OpenStudy (anonymous):

@iGreen Sorry OpenStudy crashed

OpenStudy (anonymous):

Is y 0 @iGreen

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

When x is 0, what is y?

OpenStudy (anonymous):

@sleepyjess Do you know?Because iGreen cant stay long enough...If x=0 then y would be....But if x=5 y would be 0???

OpenStudy (sleepyjess):

What?

OpenStudy (anonymous):

Read..........Tell me what you think.

OpenStudy (igreen):

??

OpenStudy (sleepyjess):

If x is 5, then y would be 32+ 5

OpenStudy (igreen):

OpenStudy just lagged out

OpenStudy (anonymous):

x=5 and y-37

OpenStudy (igreen):

Look at the table..what is 'y' when x is 0? Hours Open Number of Customers In the Store 0 12 1 17 2 22 3 27 4 32

OpenStudy (anonymous):

12

OpenStudy (igreen):

Correct, that is the y-intercept, so our final equation is: y = 5x + 12

OpenStudy (igreen):

Is that everything?

OpenStudy (anonymous):

Thank you.

OpenStudy (igreen):

No problem, Happy to help \(\Huge\ddot\smile\)